Circular No. 5960

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PU VULPECULAE H. Nussbaumer and M. Vogel, Institute of Astronomy, Zurich, write:

“International Ultraviolet Explorer observations of the symbiotic nova PU Vul show that this object has entered an eclipse (V = 10–11). This is the first eclipse after the fading in 1980, the cause of which is still debated. The IUE short-wave continuum has strongly decreased; this is due to the eclipse of the hot radiation source by the red giant. The broad emission features He II 164.0-nm, N V 124.0-nm, and N IV 171.8-nm (from the wind of the hot star) have disappeared, whereas the nebular emission lines are still present. We estimate that the end of eclipse should occur between the end of March and end of August 1994. We shall cover this period with IUE observations, but optical spectroscopy over this period is strongly desirable. Changes in visual magnitudes may remain small. However, the broad stellar wind lines should show a big increase on the way out of eclipse, in particular He II 468.6-nm, N IV 405.8-nm, and N III 464.3-nm, as well as the broad feet of the Balmer lines. Before going into eclipse, the width of these lines indicated wind velocities of ∼ 500 km/s.”

SUPERNOVA 1992bu IN NGC 3690 D. Van Buren, T. Jarrett, S. Terebey, and C. Beichman, Infrared Pro-

cessing and Analysis Center, California Institute of Technology and Jet Propulsion Laboratory; and M. Shure and C. Kaminski, NASA Infrared Telescope Facility, Institute for Astronomy, University of Hawaii, report their discovery of an apparent supernova in NGC 3690 (= Markarian 171) at α = 11^{h}25^{m}42^{s}.0, δ = +58^{o}50′10′′ (equinox 1950.0) on archival images taken in the K (2.2-μm) band as part of a search for supernovae in starburst galaxies. Offsets from Core B1 (cf. Wynn-Williams et al. 1991, Ap.J. 377, 426) of the galaxy are 5′′ east and 3′′ south. The object appears in ProtoCam data (obtained at the IRTF, Mauna Kea) for 1992 Mar. 9 UT (when K = 16.6 ± 0.2), Apr. 10 (17.2 ± 0.3), and May 6 (18.1 ± 0.5). It does not appear to limiting magnitude K ∼ 19 in NSFCam IRTF images from 1993 Dec. 27.

PERIODIC COMET KUSHIDA (1994a) Total V CCD magnitudes by H. Mikuž, Ljubljana, Slovenia (0.20-m

reflector): Mar. 4.85 UT, 11.5; 29.86, 13.1.

1994 March 30 Daniel W. E. Green